The degree measure of the vertex angle is (7x-27) the degree measure for each base angle is ( 9x-34) what is the value of one base angle
Found 2 solutions by josgarithmetic, ankor@dixie-net.com:
Answer by josgarithmetic(39632)
Answer by ankor@dixie-net.com(22740)

The degree measure of the vertex angle is (7x-27) the degree measure for each base angle is ( 9x-34) what is the value of one base angle
:
An isosceles Triangle
(7x-27) + 2(9x-34) = 180
7x - 27 + 18x - 68 = 180
25x - 95 = 180
25x = 180 + 95
25x = 275
x = 275/25
x = 11
"what is the value of one base angle?"
9(11) - 34 = 65 degrees
:
:
Check this find the vertex angle
7(11) - 27 = 50 degrees
50 + 65 + 65 = 180